OpenStudy (anonymous):

Find the center and radius of (x + 9)2 + (y + 6)2 = 9. plz help

OpenStudy (shamim):

do u mean (x+9)^2+(y+6)^2=9

OpenStudy (shamim):

u can rewrite it as (x-(-9))^2+(y-(-6))^2=3^2

OpenStudy (anonymous):

ok

OpenStudy (shamim):

now equation of a circle is (x-h)^2+(y-k)^2=a^2

OpenStudy (shamim):

now compare the above two equations nd find out the coordinates of center nd radius of a circle

OpenStudy (shamim):

plz tell me ur result

OpenStudy (mathmale):

\[(x-h)^2+(y-k)^2=a^2\] is the standard equation of a circle with center at (h,k) and radius a. If at all possible, @heyboyslol, please use Equation Editor to present your math statements for the great additional clarity it provides. Compare your (x + 9)2 + (y + 6)2 = 9 to this standard equation. Identify the values of h, h and a. With this info, write out the center, (h,k), and the radius, a.

OpenStudy (shamim):

h=-9 k=-6 now tell me a=?

OpenStudy (anonymous):

? im not really the best @ math

OpenStudy (shamim):

i think u understood nothing frm our discusion now. is not it?